# Exergetic Port-Hamiltonian Systems Modeling Language

### v1

@@ 3.1 Interconnection patterns
Figure 4: Interconnection pattern of an isothermal oscillator model. The box $\mathtt{pkc}$ represents the coupling of the potential energy domain represented by the junction on its left and the kinetic energy domain represented by the junction on its right. The box $\mathtt{env}$ represents the isothermal environment which absorbs the heat that is generated by the damping. The port $\mathtt{p}$ of the outer box exposes the kinetic energy domain, allowing for an external forcing of the oscillator model.
Figure 4 shows an interconnection pattern. The mathematical content of such a diagram is captured by the following definition:
@@
    A package of interfaces $\beta\in\mathbf{Pkg}(\mathbf{Interface})$. Names in $\operatorname{dom}(\beta)$ refer to inner boxes, $\sum\beta$ is called inner interface and names in $\operatorname{dom}(\sum\beta)$ refer to inner ports.
@@
    An interface $I$ called outer interface with ports called outer ports.
@@
      and all ports have the same associated quantity. ∎
Although the names of inner ports and outer ports are distinguished by a prefix, usually the namespaces of $\sum\beta$ and $I$ are already disjoint. We then abuse notation and simply write for instance $\mathtt{p}$ rather than $\mathtt{outer.p}$.
We graphically depict interconnection patterns as follows. Each element $b\in\operatorname{dom}(\beta)$ is the name of an inner box, which we draw as a circle. The ports of its interface $\beta(b)$ are drawn as lines emanating from the box. The ports of the outer interface $I$ are drawn as lines to the surrounding outer box. Each subset $j\in J$ corresponds to a junction drawn as a black dot. The elements of $j\subseteq P$ are precisely the connected ports. To avoid confusion, we recall that the identifier isothermal_osc written on the top left of the outer box refers to the entire pattern (akin to a function name), rather than just to the outer box (which similar to the return value of a function does not need a name).
@@
Figure 5: Interconnection pattern of an isothermal oscillator model. Whenever all ports connected to a certain junction have the same name we write the name only once at the junction.
@@
$\operatorname{dom}(\beta)\>=\>\{\mathtt{spring},\,\mathtt{pkc},\,\mathtt{mass},\,\mathtt{damping},\,\mathtt{env}\}\,.$
The function $\beta:\operatorname{dom}(\beta)\rightarrow\mathbf{Interface}$ is given by
$\begin{split}\mathtt{spring}\>&\mapsto\>(\{\mathtt{q}\},\,\tau_{\text{spring}})\\ \mathtt{pkc}\>&\mapsto\>(\{\mathtt{q},\,\mathtt{p}\},\,\tau_{\text{pkc}})\\ \mathtt{mass}\>&\mapsto\>(\{\mathtt{p}\},\,\tau_{\text{mass}})\\ \mathtt{damping}\>&\mapsto\>(\{\mathtt{p},\,\mathtt{s}\},\,\tau_{\text{damping}})\\ \mathtt{env}\>&\mapsto\>(\{\mathtt{s}\},\,\tau_{\text{env}})\end{split}$
with $\tau_{\text{spring}}(\mathtt{q})=((\mathbb{R},\,\mathtt{displacement}),\,\mathtt{p})$, etc. Hence, the namespace of the inner interface $\sum\beta$ is
$\operatorname{dom}(\textstyle\sum\beta)\>=\>\{\mathtt{spring.q},\,\mathtt{pkc.q},\,\mathtt{pkc.p},\,\mathtt{mass.p},\,\mathtt{damping.p},\,\mathtt{damping.s},\,\mathtt{env.s}\}\,.$
@@
$\bigl{\{}\{\mathtt{spring.q},\,\mathtt{pkc.q}\},\,\{\mathtt{mass.p},\,\mathtt{pkc.p},\,\mathtt{damping.p},\,\mathtt{p}\},\,\{\mathtt{env.s},\,\mathtt{damping.s}\}\bigr{\}}\,.\qed$
@@
[…]
In this we implicitly let $\operatorname{dom}(\beta)=\{\mathtt{I_{1}},\,\ldots,\,\mathtt{I_{n}}\}$ and $\beta(\mathtt{I_{i}})=I_{i}$. The left arrow is a surjection because each junction has at least one connected inner port and the right arrow is an injection because each junction has at most one connected outer port. Interconnection patterns are a special case of undirected wiring diagrams Spivak (2013), which do not have the injectivity and surjectivity conditions.
As shown in Figure 1, interconnection patterns may have another feature that is not formalized by 3.1, namely port combinators for combining ports into multiports. Although namespaces seem to provide an adequate basis to formalize multiports and port combinators, we defer this to later work. At this point, we simply regard them as ad-hoc syntactic sugar for expressing parallel connections in a tidy and efficient manner.
@@ 3.2 Composition of interconnection patterns
Whenever the outer interface of one pattern and the interface corresponding to an inner box of another pattern are equivalent up to renaming of the ports, the former pattern can be substituted into the latter, with no further data required. This is called composition. Having a compositional syntax facilitates dealing with complex systems, as it allows to encapsulate subsystems, which can be easily reused and replaced, as long as interfaces match. As an example, the isothermal oscillator model can be refined into a nonisothermal oscillator model by substituting a more detailed damper model.
Figure 6: Interconnection pattern of a nonisothermal damper model. The box $\mathtt{mf}$ represents a mechanical friction model connected to the kinetic energy domain represented by the outer port $\mathtt{p}$. The box $\mathtt{tc}$ represents a thermal capacity, which can store the heat generated by friction. The box $\mathtt{ht}$ represents a heat transfer model, which allows the thermal capacity to reach equilibrium with the thermal energy domain represented by the outer port $\mathtt{s}$.
Figure 7: Interconnection pattern of a nonisothermal oscillator model.
Substituting the pattern in Figure 6 into the pattern in Figure 5, yields the pattern in Figure 7. Graphically, the former pattern is inserted into the inner box called $\mathtt{damping}$ of the latter pattern. In this step, the named sum for interfaces is used. Then, the common interface is deleted and for each of its ports, the two assigned junctions are identified.
Category theory provides a language to talk about composition and it includes well-behavedness conditions, which ensure that composition works in the way we expect, just as we need not worry about the difference between $(1+2)+1$ and $1+(2+1)$. In a following paper, we will formalize this using the language of operads.
@@ 4 Semantics of interconnection patterns
In this section, we discuss how to interpret an interconnection pattern as imposing a relation between the port variables of its inner and outer interface. We do this in two parts. In the first part, we talk about the relation as a collection of equations and discuss its physical interpretation. In the second part, we reify it as a geometric object, namely as a relation between bundles of port variables.
@@ 4.1 Interconnection patterns as equations
Informally, the semantics of an interconnection pattern $(\beta,\,I,\,J)$ can be seen as a collection of equations. Specifically, for each junction $j\in J$, we have the following:
@@
As detailed in Section 7, combining these equations with equations that determine the semantics of each of the subsystems filling the inner boxes yields a collection of equations that determines the semantics of the resulting composite system.
Since all ports connected to a junction have the same quantity, their associated state variables take values in the same space, making equality of state a well-typed equation. As a consequence of equality of state, all effort variables take values in the same fiber of the same cotangent bundle, making equality of effort also well-typed. Similarly, all flow variables take values in the same fiber of the same tangent bundle, making equality of net flow well-typed.
Also note that in the equality of net flow, the inner ports are on the left hand side and the outer ports are on the right hand side. As a consequence, at every junction $j\in J$ the following power balance equation is satisfied:
@@
Summing over all junctions on both sides, we get that the net power supplied to all subsystems is equal to the net power supplied to the composite system. Taking into account also the choice of a sign made in Sections 6.1 and 6.2, we see that this agrees with the sign convention used in thermodynamics, stating that power supplied to any system, be it a subsystem or the composite system, is counted with a positive sign.
Junctions correspond to energy domains. Hence, an outer port exposes an energy domain of the respective composite system. Since this interpretation also applies to the primitive systems discussed in Section 6, we can say that ports also represent energy domains and whenever two ports are connected to the same junction the energy domains are identified. In other words, interconnection means that systems share common energy domains. The state of a shared energy domain is given by the shared state variable, according to the equality of state. The connected systems may exchange the respective physical quantity according to the equality of net flow. Moreover, such an exchange also implies an exchange of energy, according to the equality of effort.
@@
Let’s again consider the pattern in Figure 7. At the junction representing the kinetic energy domain, the momentum variable (state) is shared and forces (flows) balance at equal velocity (effort):
@@
We can now briefly reconsider the surjectivity and injectivity conditions which distinguish interconnection patterns form undirected wiring diagrams. The surjectivity condition says that each junction must have at least one connected inner port. In other words, an energy domain of the composite system must belong to at least one of its subsystems. Of course, an energy domain that is not part of the system cannot be exposed. The injectivity condition says that each junction may have at most one connected outer port. In other words, an energy domain can be exposed at most once. Exposing an energy domain twice and identifying it again at a higher level of nesting by connecting the two respective ports would lead to ill-defined flow variables.
@@ 4.2 Interconnection patterns as relations
Similarly to this prototypical example, the relation assigned to an interconnection pattern is given by a subbundle of the Cartesian product of the two bundles of port variables associated to the inner and the outer interface.
@@
Let $P=(\beta,\,I,\,J)$ be an interconnection pattern and let $I_{\text{in}}=\sum\beta$ denote its inner interface. Further, let $\mathcal{P}_{I_{\text{in}}}\twoheadrightarrow\mathcal{X}_{I_{\text{in}}}$ be the bundle of port variables associated to the inner interface and let $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ be the bundle of port variables associated to the outer interface. Based on the partition $J$, let $\mathcal{P}_{P}\subset\mathcal{P}_{I_{\text{in}}}\times\mathcal{P}_{I}$ be the subspace where equality of state, equality of effort, and equality of net flow are satisfied. Similarly, let $\mathcal{X}_{P}\subset\mathcal{X}_{I_{i}}\times\mathcal{X}_{I}$ be the subspace where equality of state is satisfied. Then, the semantics assigned to interconnection pattern $P$ is the relation $\mathcal{P}_{P}:\mathcal{P}_{I_{\text{in}}}\nrightarrow\mathcal{P}_{I}$ between the port variables of the inner interface and the port variables of the outer interface that is given by the vector subbundle $\mathcal{P}_{P}\twoheadrightarrow\mathcal{X}_{P}$. ∎
It can be shown that this assignment of semantics is functorial, which means that it makes no difference whether one first composes some interconnection patterns and then asks for the relation assigned to the composite or one first asks for the relations assigned to the original patterns and then composes these relations.
Figure 8: Interconnection pattern of a DC shunt motor model. The stator and rotor are connected to a common electric energy domain that is exposed via the outer port $\mathtt{q}$. The kinetic energy domain of the rotor is exposed via the outer port $\mathtt{p}$. The conversion between electric and kinetic energy in the rotor depends on the magnetic energy domain of the stator, which is represented by the junction with connected state ports named $\mathtt{b_{k}}$. The full model is discussed in Section 9.
@@
Consider the interconnection pattern $P=(\beta,\,I,\,J)$ depicted in Figure 8 with names of inner boxes given by $\operatorname{dom}(\beta)=\{\mathtt{stator},\,\mathtt{rotor}\}$ and interfaces given by
$\beta(\mathtt{stator})\>=\>\left(\{\mathtt{q},\,\mathtt{b_{s}}\},\,\tau_{\text{s}}\right)$
$\beta(\mathtt{rotor})\>=\>\left(\{\mathtt{q},\,\mathtt{p},\,\mathtt{b_{s}}\},\,\tau_{\text{r}}\right)$
$I\>=\>\left(\{\mathtt{q},\,\mathtt{p}\},\,\tau\right)\,,$
@@
$\tau_{\text{s}}(\mathtt{q})\>=\>\tau_{\text{r}}(\mathtt{q})\>=\>\tau(\mathtt{q})\>=\>((\mathbb{R},\,\mathtt{charge}),\,\mathsf{p})$
$\tau_{\text{s}}(\mathtt{b_{s}})\>=\>\tau_{\text{r}}(\mathtt{b_{s}})\>=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{s})$
$\tau_{\text{r}}(\mathtt{p})\>=\>\tau(\mathtt{p})\>=\>((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})\,.$
Furthermore, the interconnection is determined by
$J\>=\>\bigl{\{}\{\mathtt{stator.q},\,\mathtt{rotor.q},\,\mathtt{q}\},\,\{\mathtt{stator.b_{s}},\,\mathtt{rotor.b_{s}}\},\,\{\mathtt{rotor.p},\,\mathtt{p}\}\bigr{\}}\,.$
@@
$\begin{split}\mathcal{P}_{I_{\text{in}}}\>=\>&\mathbb{T}\mathbb{R}\times\mathbb{R}\times\mathbb{T}\mathbb{R}\times\mathbb{T}\mathbb{R}\times\mathbb{R}\\ \>\ni\>&\bigl{(}(\mathtt{stator.q.x},\,\mathtt{stator.q.f},\,\mathtt{stator.q.e}),\,\mathtt{stator.b_{s}.x},\,\\ &\>\>(\mathtt{rotor.q.x},\,\mathtt{rotor.q.f},\,\mathtt{rotor.q.e}),\,\\ &\>\>(\mathtt{rotor.p.x},\,\mathtt{rotor.p.f},\,\mathtt{rotor.p.e}),\,\mathtt{rotor.b_{s}.x}\bigr{)}\end{split}$
@@
$\mathcal{P}_{I}\>=\>\mathbb{T}\mathbb{R}\times\mathbb{T}\mathbb{R}\>\ni\>\bigl{(}(\mathtt{q.x},\,\mathtt{q.f},\,\mathtt{q.e}),\,(\mathtt{p.x},\,\mathtt{p.f},\,\mathtt{p.e})\bigr{)}\,.$
@@
$\begin{split}\mathcal{P}_{P}\>=\>\bigl{\{}\bigl{(}&(\mathtt{stator.q.x},\,\mathtt{stator.q.f},\,\mathtt{stator.q.e})\,,\,\mathtt{stator.b_{s}.x}\,,\,\\ &(\mathtt{rotor.q.x},\,\mathtt{rotor.q.f},\,\mathtt{rotor.q.e})\,,\,\\ &(\mathtt{rotor.p.x},\,\mathtt{rotor.p.f},\,\mathtt{rotor.p.e})\,,\,\mathtt{rotor.b_{s}.x}\,,\,\\ &(\mathtt{q.x},\,\mathtt{q.f},\,\mathtt{q.e})\,,\,(\mathtt{p.x},\,\mathtt{p.f},\,\mathtt{p.e})\bigr{)}\in\mathcal{P}_{I_{i}}\times\mathcal{P}_{I}\,\mid\,\\ &\mathtt{stator.q.x}=\mathtt{rotor.q.x}=\mathtt{q.x}\,,\,\\ &\mathtt{stator.q.e}=\mathtt{rotor.q.e}=\mathtt{q.e}\,,\,\\ &\mathtt{stator.q.f}+\mathtt{rotor.q.f}=\mathtt{q.f}\,,\,\\ &\mathtt{stator.b_{s}.x}=\mathtt{rotor.b_{s}.x}\,,\,\\ &\mathtt{rotor.p.x}=\mathtt{p.x}\,,\,\\ &\mathtt{rotor.p.e}=\mathtt{p.e}\,,\,\\ &\mathtt{rotor.p.f}=\mathtt{p.f}\,\bigr{\}}\,.\qed\end{split}$

### v2

@@ 3.1 Interconnection patterns
Figure 4: Interconnection pattern of an isothermal oscillator model. Box $\mathtt{pkc}$ represents the reversible coupling of the potential energy domain represented by the junction on its left and the kinetic energy domain represented by the junction on its right. Box $\mathtt{env}$ represents the isothermal environment, which absorbs the heat that is generated by the damping. The outer box represents the interface of the composite system. Its port $\mathtt{p}$ exposes the kinetic energy domain, allowing for an external forcing of the oscillator model.
Figure 4 shows an interconnection pattern. The mathematical content of such a diagram is captured by the following definition.
@@
    A directory of interfaces $\beta\in\mathsf{Dtry}(\mathsf{Interface})$. Names in $\operatorname{dom}(\beta)$ refer to inner boxes, $\sum\beta$ is called inner interface and names in $\operatorname{dom}(\sum\beta)$ refer to inner ports.
@@
    An interface $I$ called outer interface, whose ports are called outer ports.
@@
      and all ports in $j$ have the same associated quantity. ∎
To prevent name conflicts, the definition distinguishes inner ports and outer ports with the prefixes $\mathtt{inner}$ and $\mathtt{outer}$. While preventing name conflicts is crucial for implementation of the modeling language, we omit these prefixes for conciseness in what follows.
We graphically depict interconnection patterns as follows. Each element $b\in\operatorname{dom}(\beta)$ is the name of an inner box, which we draw as a circle. The ports of its interface $\beta(b)$ are drawn as lines emanating from the box. The ports of the outer interface $I$ are drawn as lines to the surrounding outer box. Each subset $j\in J$ corresponds to a junction drawn as a black dot. The elements of $j\subseteq P$ are precisely the connected ports. To avoid confusion, we recall that the identifier isothermal_osc written on the top left of the outer box refers to the entire pattern (akin to a function name in programming), rather than just to the outer box (which similar to the return value of a function does not need a name).
@@
Figure 5: Interconnection pattern of an isothermal oscillator model. Whenever all ports connected to a certain junction have the same name, we write the name only once at the junction.
@@
$\operatorname{dom}(\beta)\>=\>\{\mathtt{damping},\,\mathtt{env},\,\mathtt{mass},\,\mathtt{pkc},\,\mathtt{spring}\}\,.$
The function $\beta:\operatorname{dom}(\beta)\rightarrow\mathsf{Interface}$ is given by
$\begin{split}\mathtt{damping}\>&\mapsto\>(\{\mathtt{p},\,\mathtt{s}\},\,\tau_{\text{damping}})\\ \mathtt{env}\>&\mapsto\>(\{\mathtt{s}\},\,\tau_{\text{env}})\\ \mathtt{mass}\>&\mapsto\>(\{\mathtt{p}\},\,\tau_{\text{mass}})\\ \mathtt{pkc}\>&\mapsto\>(\{\mathtt{p},\,\mathtt{q}\},\,\tau_{\text{pkc}})\\ \mathtt{spring}\>&\mapsto\>(\{\mathtt{q}\},\,\tau_{\text{spring}})\end{split}$
with $\tau_{\text{damping}}(\mathtt{p})=((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p})$, etc. Hence, the namespace of the inner interface $\sum\beta$ is
$\operatorname{dom}(\textstyle\sum\beta)\>=\>\{\mathtt{damping.p},\,\mathtt{damping.s},\,\mathtt{env.s},\,\mathtt{mass.p},\,\mathtt{pkc.p},\,\mathtt{pkc.q},\,\mathtt{spring.q}\}\,.$
@@
$\bigl{\{}\{\mathtt{damping.p},\,\mathtt{mass.p},\,\mathtt{pkc.p},\,\mathtt{p}\},\,\{\mathtt{pkc.q},\,\mathtt{spring.q}\},\,\{\mathtt{damping.s},\,\mathtt{env.s}\}\bigr{\}}\,.\qed$
@@
${\sum\left[\mathtt{I_{1}}\mapsto I_{1},\,\ldots,\,\mathtt{I_{n}}\mapsto I_{n}\right]}$${J}$${I.}$
In this we implicitly let $\operatorname{dom}(\beta)=\{\mathtt{I_{1}},\,\ldots,\,\mathtt{I_{n}}\}$ and $\beta(\mathtt{I_{i}})=I_{i}$. The left arrow is a surjection because each junction has at least one connected inner port and the right arrow is an injection because each junction has at most one connected outer port. Interconnection patterns are a special case of undirected wiring diagrams [16], which do not have the injectivity and surjectivity conditions.
Multiports are just multiple ports with a common prefix. As demonstrated in Figure 2, this enables a syntactic sugar for expressing parallel connections in a tidy and efficient manner. We defer the details to later work, as the examples considered here make no use of this.
@@ 3.2 Composition of interconnection patterns
Whenever the outer interface of one pattern and the interface corresponding to an inner box of another pattern are equal (up to a given renaming of the ports), the former pattern can be substituted into the latter, with no further data required. This is called composition. Having a compositional syntax facilitates dealing with complex systems, as it enables the encapsulation of subsystems. These can be easily reused and replaced, as long as interfaces match. As an example, the isothermal oscillator model can be refined into a nonisothermal oscillator model based on the damper model in Figure 6.
Figure 6: Interconnection pattern of a nonisothermal damper model. Box $\mathtt{mf}$ represents a mechanical friction model connected to the kinetic energy domain represented by the outer port $\mathtt{p}$. Box $\mathtt{tc}$ represents a thermal capacity, which stores the heat generated by friction. Box $\mathtt{ht}$ represents a heat transfer model, which allows the thermal capacity to reach equilibrium with the thermal energy domain represented by the outer port $\mathtt{s}$.
Figure 7: Interconnection pattern of a nonisothermal oscillator model. Except for the graphical layout, the pattern is uniquely defined as the composite of the patterns in Figures 5 and 6.
Substituting the pattern in Figure 6 into the pattern in Figure 5, yields the pattern in Figure 7. Graphically, the former pattern is inserted into the inner box called $\mathtt{damping}$ of the latter pattern. The common interface is deleted and for each of its ports, the two assigned junctions (one in each pattern) are identified. The named sum for interfaces is used to obtain the inner interface of the composed pattern.
Category theory provides a language to talk about composition, and it includes well-behavedness conditions, which ensure that composition works in the way we expect, just as we need not worry about the difference between $(1+2)+3$ and $1+(2+3)$. A future paper shall formalize the composition of interconnection patterns based on directories and multicategories/operads. To limit the scope of this article, we just mention that it is not difficult to see why the procedure for composition outlined in the previous paragraph preserves the three conditions in 3.1.
@@ 4 Semantics of interconnection patterns
In Section 2.5 we have defined bundles of port variables associated to interfaces. In Section 3.1 we have defined interconnection patterns, which have an inner interface (containing all ports of the subsystems) and an outer interface (containing the ports of the composite system). We now discuss how to interpret an interconnection pattern as imposing a relation between the port variables of its inner interface and the port variables of its outer interface. We do this in two parts. In the first part, we talk about the relation as a collection of equations, and we discuss its physical interpretation. In the second part, we reify it as a geometric object, namely as a relation between bundles of port variables.
@@ 4.1 Interconnection patterns as equations
The semantics of an interconnection pattern $(\beta,\,I,\,J)$ can be seen as a collection of equations. Specifically, for each junction $j\in J$, we have the following.
@@
We note that the sum on the left-hand side has at least one term (surjectivity condition) and the sum on the right-hand side has at most one term (injectivity condition).
As detailed in Section 7, combining these equations with equations that determine the semantics of the subsystems filling the inner boxes yields a collection of equations that determines the semantics of the resulting composite system.
Since all ports connected to a junction have the same quantity, their associated state variables take values in the same space, making equality of state a well-typed equation. As a consequence of this, all effort variables take values in the same fiber of the same cotangent bundle, making equality of effort also well-typed. Similarly, all flow variables take values in the same fiber of the same tangent bundle, making equality of net flow well-typed.
Equality of state is required, since the behavior of one system may in general depend on the state of another system, even if the two systems do not exchange energy. We refer to Figures 8 and 9 for an example of this.
As a consequence of equality of effort and equality of net flow, at every junction $j\in J$ the following power balance equation is satisfied:
@@
Summing over all junctions on both sides gives that the net power supplied to all subsystems is equal to the net power supplied to the composite system.
Junctions correspond to energy domains and so do (outer) ports, since they expose energy domains. Whenever two ports are connected to the same junction the respective energy domains are identified. This is exactly how interconnection patterns compose. Interconnection means that systems share energy domains. The state of a shared energy domain is given by the shared state variable, according to the equality of state. The connected systems may exchange the respective physical quantity according to the equality of net flow. Moreover, based on the equality of effort, this exchange implies a transfer of (exergetic) power.
@@
Let’s again consider the pattern in Figure 7. At the junction which represents the kinetic energy domain, the momentum variable (state) is shared and forces (flows) balance at equal velocity (effort):
@@
We can now make sense of the surjectivity and injectivity conditions distinguishing interconnection patterns form undirected wiring diagrams. The surjectivity condition says that each junction must have at least one connected inner port. In other words, an energy domain can only belong to a composite system if it belongs to at least one subsystem. The injectivity condition says that each junction may have at most one connected outer port. In other words, an energy domain can be exposed at most once. Exposing an energy domain twice and identifying it again at a higher level by connecting the two respective ports would lead to undefined flow variables.
@@ 4.2 Interconnection patterns as relations
Analogous to the prototypical example of a relation between two sets, the relation assigned to an interconnection pattern is defined by a vector subbundle of the Cartesian product of the bundle of port variables associated to the inner interface and the bundle of port variables associated to the outer interface.
@@
Let $P=(\beta,\,I,\,J)$ be an interconnection pattern and let $I_{\text{in}}=\sum\beta$ denote its inner interface. Further, let $\mathcal{P}_{I_{\text{in}}}\twoheadrightarrow\mathcal{X}_{I_{\text{in}}}$ be the bundle of port variables associated to the inner interface, and let $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ be the bundle of port variables associated to the outer interface. Based on the partition $J$, let $\mathcal{P}_{P}\subset\mathcal{P}_{I_{\text{in}}}\times\mathcal{P}_{I}$ be the subspace where equality of state, equality of effort, and equality of net flow are satisfied. Similarly, let $\mathcal{X}_{P}\subset\mathcal{X}_{I_{i}}\times\mathcal{X}_{I}$ be the subspace where equality of state is satisfied. Then, the semantics assigned to interconnection pattern $P$ is the relation $\mathcal{P}_{P}:\mathcal{P}_{I_{\text{in}}}\nrightarrow\mathcal{P}_{I}$ between the port variables of the inner interface and the port variables of the outer interface that is defined by the vector subbundle $\mathcal{P}_{P}\twoheadrightarrow\mathcal{X}_{P}$. ∎
Figure 8: Interconnection pattern of the DC shunt motor model discussed in Section 9. The stator and rotor are connected to a common electric energy domain that is exposed via the outer port $\mathtt{q}$. The kinetic energy domain of the rotor is exposed via the outer port $\mathtt{p}$. The conversion between electric and kinetic energy in the rotor depends on the magnetic energy domain of the stator, which is represented by the junction with two connected state ports named $\mathtt{b_{s}}$.
@@
Consider the interconnection pattern $P=(\beta,\,I,\,J)$ depicted in Figure 8 with names of inner boxes given by $\operatorname{dom}(\beta)=\{\mathtt{rotor},\,\mathtt{stator}\}$ and interfaces given by
$\beta(\mathtt{rotor})\>=\>\left(\{\mathtt{b_{s}},\,\mathtt{p},\,\mathtt{q}\},\,\tau_{\text{r}}\right)$
$\beta(\mathtt{stator})\>=\>\left(\{\mathtt{b_{s}},\,\mathtt{q}\},\,\tau_{\text{s}}\right)$
$I\>=\>\left(\{\mathtt{p},\,\mathtt{q}\},\,\tau\right)\,,$
@@
$\tau_{\text{r}}(\mathtt{b_{s}})\>=\>\tau_{\text{s}}(\mathtt{b_{s}})\>=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{s})$
$\tau_{\text{r}}(\mathtt{p})\>=\>\tau(\mathtt{p})\>=\>((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$
$\tau_{\text{r}}(\mathtt{q})\>=\>\tau_{\text{s}}(\mathtt{q})\>=\>\tau(\mathtt{q})\>=\>((\mathbb{R},\,\mathtt{charge}),\,\mathsf{p})\,.$
Further, the interconnection is determined by
$J\>=\>\bigl{\{}\{\mathtt{rotor.b_{s}},\,\mathtt{stator.b_{s}}\},\,\{\mathtt{rotor.p},\,\mathtt{p}\},\,\{\mathtt{rotor.q},\,\mathtt{stator.q},\,\mathtt{q}\}\bigr{\}}\,.$
@@
$\begin{split}\mathcal{P}_{I_{\text{in}}}\>=\>&\mathbb{R}\times\mathbb{T}\mathbb{R}\times\mathbb{T}\mathbb{R}\times\mathbb{R}\times\mathbb{T}\mathbb{R}\\ \>\ni\>&\bigl{(}\mathtt{rotor.b_{s}.x},\,(\mathtt{rotor.p.x},\,\mathtt{rotor.p.f},\,\mathtt{rotor.p.e}),\,\\ &\>\>(\mathtt{rotor.q.x},\,\mathtt{rotor.q.f},\,\mathtt{rotor.q.e}),\,\\ &\>\>\mathtt{stator.b_{s}.x},\,(\mathtt{stator.q.x},\,\mathtt{stator.q.f},\,\mathtt{stator.q.e})\bigr{)}\end{split}$
@@
$\mathcal{P}_{I}\>=\>\mathbb{T}\mathbb{R}\times\mathbb{T}\mathbb{R}\>\ni\>\bigl{(}(\mathtt{p.x},\,\mathtt{p.f},\,\mathtt{p.e}),\,(\mathtt{q.x},\,\mathtt{q.f},\,\mathtt{q.e})\bigr{)}\,.$
@@
$\begin{split}\mathcal{P}_{P}\>=\>\bigl{\{}&\bigl{(}\mathtt{rotor.b_{s}.x}\,,\,(\mathtt{rotor.p.x},\,\mathtt{rotor.p.f},\,\mathtt{rotor.p.e})\,,\,\\ &(\mathtt{rotor.q.x},\,\mathtt{rotor.q.f},\,\mathtt{rotor.q.e})\,,\,\\ &\mathtt{stator.b_{s}.x}\,,\,(\mathtt{stator.q.x},\,\mathtt{stator.q.f},\,\mathtt{stator.q.e})\,,\,\\ &(\mathtt{p.x},\,\mathtt{p.f},\,\mathtt{p.e})\,,\,(\mathtt{q.x},\,\mathtt{q.f},\,\mathtt{q.e})\bigr{)}\in\mathcal{P}_{I_{i}}\times\mathcal{P}_{I}\,\mid\,\\ &\mathtt{rotor.b_{s}.x}=\mathtt{stator.b_{s}.x}\,,\,\\ &\mathtt{rotor.p.x}=\mathtt{p.x}\,,\,\\ &\mathtt{rotor.p.f}=\mathtt{p.f}\,,\,\\ &\mathtt{rotor.p.e}=\mathtt{p.e}\,,\,\\ &\mathtt{rotor.q.x}=\mathtt{stator.q.x}=\mathtt{q.x}\,,\,\\ &\mathtt{rotor.q.f}+\mathtt{stator.q.f}=\mathtt{q.f}\,,\,\\ &\mathtt{rotor.q.e}=\mathtt{stator.q.e}=\mathtt{q.e}\,\bigr{\}}\,.\qed\end{split}$
Given a hierarchy of interconnection patterns, we can first compose the patterns and then determine the relation assigned to the composed pattern. Alternatively, we can first determine the relation assigned to each individual pattern and then compose these relations. Both ways lead to the same result. The combinatorial syntax and its relational semantics hence compose in compatible ways. In the language of category theory, we say that the (assignment of) semantics is functorial. A future paper shall use directories and the theory of multicategories/operads to demonstrate that 4.2 indeed defines a suitable notion of functor.